# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ VI-A Proof of Theorem 4
We will denote the distribution corresponding to the pairwise comparison matrix $P_{0}$ by ${\mathbb{P}}_{0}$. Additionally, note that under $H_{0}$, the stationary distribution of the canonical Markov matrix ${S}$ is uniform, i.e., $\pi=\frac{1}{n}\mathbf{1}_{n}$. Under the alternative hypothesis, we assume that the pairwise comparison matrix $P_{\theta}$ is generated by sampling the parameter $\theta$ uniformly from the set $\Theta$, i.e.,
@@
where $\Theta$ is set of all permuation matrices, and ${Q}_{\theta}$ is the $\frac{n}{2}\times\frac{n}{2}$ permutation matrix corresponding to the permutation $\theta$ and the perturbation $\eta\in(0,\frac{1}{2})$. Let ${\mathbb{P}}_{\Theta}$ denote the overall mixture distribution and ${\mathbb{P}}_{\theta}$ denotes the distribution corresponding to the pairwise comparison matrix $P_{\theta}$. The construction of this mixture was inspired by [49]. However, there are two notable differences. Firstly, the problem in [49] is distinguishing whether two sets of data samples consisting of pairwise comparisons are coming from the same underlying distribution or two different distributions described by a pairwise comparison model. In contrast, our work tests whether or not a single dataset is sampled from a BTL model. Secondly, the manner in which a notion of distance is used to define the deviation of the given data from the null hypothesis is different in the two works.
@@
$\|\Pi P_{\theta}+P_{\theta}^{\mathrm{T}}\Pi-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}^{2}$ $=\frac{n}{2}\left((x+y)\left(\frac{1}{2}+\eta\right)-y\right)^{2}+\frac{n}{2}\left((x+y)\left(\frac{1}{2}-\eta\right)-x\right)^{2}$ (89)
@@
Let $\epsilon=\|\Pi P_{\theta}+P_{\theta}\Pi-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}/(n\|\pi\|_{\infty})$ to ensure that the $P_{\theta}$’s satisfy the condition of the alternative hypothesis in (19). Substituting the values of $\|\pi\|_{\infty}=y$ and $\|\Pi P_{\theta}+P_{\theta}^{\mathrm{T}}\Pi-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}$ implies that
$\epsilon^{2}\leq C\frac{\eta^{2}}{n},$ (90)
for some constant $C>0$.
Now, the Ingster-Suslina method [71] states that
$\mathcal{R}_{\mathsf{m}}\geq 1-\sqrt{\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})},$ (91)
where $\chi^{2}(\cdot||\cdot)$ denotes the $\chi^{2}$-divergence. We compute $\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ by expressing it as an expectation with respect to two independent pairwise models corresponding to permutation $\theta$ and $\theta^{\prime}$ drawn independently and uniformly at random from $\Theta$ as
$\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ $={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\left[\int\frac{d{\mathbb{P}}_{\theta}d{\mathbb{P}}_{\theta^{\prime}}}{d{\mathbb{P}}_{0}}\right],$
where $d{\mathbb{P}}_{\theta}$ denotes the measure induced by the pairwise comparison model corresponding to permutation $\theta$. Let $p_{ij}$ and $p_{ij}^{\prime}$ be the pairwise probabilities corresponding to permutations $\theta$ and $\theta^{\prime}$. Now we will utilize the tensorization property of $1+\chi^{2}(P||Q)$, i.e., we utilize the fact that for distributions $P_{1},Q_{1},\dots,P_{n},Q_{n}$, we have
@@
Therefore, the $\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ can be simplified as
$1+\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ $={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\left[\prod_{i=1}^{n}\prod_{\begin{subarray}{c}j=1:\\ j\neq i\end{subarray}}^{n}\bigg{(}\sum_{m=0}^{k}\frac{{k\choose m}(p_{ij})^{m}(1-p_{ij})^{k-m}{k\choose m}(p_{ij}^{\prime})^{m}(1-p^{\prime}_{ij})^{k-m}}{{k\choose m}\left(\frac{1}{2}\right)^{k}}\bigg{)}\right]$
$={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\left[\prod_{i=1}^{n}\prod_{\begin{subarray}{c}j=1:\\ j\neq i\end{subarray}}^{n}\bigg{(}\sum_{m=0}^{k}\frac{{k\choose m}(p_{ij})^{m}(1-p_{ij})^{k-m}(p_{ij}^{\prime})^{m}(1-p^{\prime}_{ij})^{k-m}}{\left(\frac{1}{2}\right)^{k}}\bigg{)}\right].$ (92)
We will focus on the $(i,j)$th term of the product in Equation 92 for $i\neq j$ and denote it as $f(p_{ij},p^{\prime}_{ij})$:
@@
By our construction of pairwise comparison matrices both $p_{ij},p^{\prime}_{ij}$ take values in set $\{\frac{1}{2},\frac{1}{2}+\eta\}$ if $j\geq i$ (and $\{\frac{1}{2},\frac{1}{2}-\eta\}$ otherwise). Furthermore, whenever either $p_{ij}$ or $p_{ij}^{\prime}$ equals $\frac{1}{2}$, we have $f(p_{ij},p^{\prime}_{ij})=1$. Additionally, by Eq. 93 we have $f(\frac{1}{2}-\eta,\frac{1}{2}-\eta)=f(\frac{1}{2}+\eta,\frac{1}{2}+\eta)$. Let a random variable $B$ denote the number of occurrences where $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}+\eta$ in randomly drawn permutation $\theta$ and $\theta^{\prime}$ (or equivalently, $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}-\eta$). Consequently, we obtain
@@
This is because by translated skew-symmetry the number of occurrences of $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}+\eta$ is the same as the occurrences where $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}-\eta$ and $f(p_{ij},p_{ij}^{\prime})$ is same in both the scenarios. Moreover, we can simplify $f(p_{ij},p_{ij}^{\prime})$ as
$f\left(\frac{1}{2}+\eta,\frac{1}{2}+\eta\right)$ $=2^{k}\sum_{m=0}^{k}{k\choose m}\left(\frac{1}{2}+\eta\right)^{2m}\left(\frac{1}{2}-\eta\right)^{2k-2m}$
@@
$=2^{k}\left(\frac{1}{2}+2\eta^{2}\right)^{k}\sum_{m=0}^{k}{k\choose m}\left(\frac{1}{2}+\frac{\eta}{\frac{1}{2}+2\eta^{2}}\right)^{m}\left(\frac{1}{2}-\frac{\eta}{\frac{1}{2}+2\eta^{2}}\right)^{k-m}$
@@
$1+\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ $={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\bigg{[}(1+4\eta^{2})^{2kB}\bigg{]}=\sum_{b=0}^{n/2}{\mathbb{P}}(B=b)(1+4\eta^{2})^{2kb}.$ (95)
To derive an upper bound for ${\mathbb{P}}(B=b)$, for a fixed a permutations $\theta$, we aim to find another permutation $\theta^{\prime}$ with exactly $b$ aligned elements. These $b$ matches can be selected in (${{n/2}\choose b}$ ways. For the remaining $n/2-b$ elements we require derangements to avoid alignment with the perturbed elements in $\theta$. Therefore, the number of such permutations is $(\frac{n}{2}-b)!\sum_{i=0}^{n/2-b}\frac{(-1)^{i}}{i!}$. Clearly, this quantity is upper bounded by $\frac{1}{2}(\frac{n}{2}-b)!$, yielding an upper bound on ${\mathbb{P}}(B=b)$ as
${\mathbb{P}}(B=b)\leq\frac{{{n/2}\choose b}\times\frac{1}{2}(\frac{n}{2}-b)!}{(n/2)!}\leq\frac{1}{2}\times\frac{1}{b!}.$
@@
$\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ $\leq\sum_{b=0}^{n/2}\frac{1}{2(b!)}\times\bigg{(}(1+4\eta^{2})^{2kb}-1\bigg{)}+\sum_{b=0}^{n/2}{\mathbb{P}}(B=b)-1$

### v2

@@ VI Proof of Lower Bound
We will denote the distribution corresponding to the pairwise comparison matrix $P_{0}$ by ${\mathbb{P}}_{0}$. Additionally, note that under $H_{0}$, the stationary distribution of the canonical Markov matrix S{S} is uniform, i.e., $\pi=\frac{1}{n}\mathbf{1}_{n}$. Under the alternative hypothesis, we assume that the pairwise comparison matrix $P_{\theta}$ is generated by sampling the parameter $\theta$ uniformly from the set $\Theta$, i.e.,
@@
where $\Theta$ is set of all permutation matrices, and ${Q}_{\theta}$ is the $\frac{n}{2}\times\frac{n}{2}$ permutation matrix corresponding to the permutation $\theta$ and the perturbation $\eta\in(0,\frac{1}{2})$. Let ${\mathbb{P}}_{\Theta}$ denote the overall mixture distribution and ${\mathbb{P}}_{\theta}$ denotes the distribution corresponding to the pairwise comparison matrix $P_{\theta}$. The construction of this mixture was inspired by [RastogiBalakrishnanShahAarti2022]. However, there are two notable differences. Firstly, the problem in [RastogiBalakrishnanShahAarti2022] is distinguishing whether two sets of data samples consisting of pairwise comparisons are coming from the same underlying distribution or two different distributions described by a pairwise comparison model. In contrast, our work tests whether or not a single dataset is sampled from a BTL model. Secondly, the manner in which a notion of distance is used to define the deviation of the given data from the null hypothesis is different in the two works.
@@
$\left\|\Pi P_{\theta}+P_{\theta}\Pi-\mathbf{1}_{n}\pi^{\mathrm{T}}\right\|_{\mathrm{F}}^{2}=$ (87)
$\ \frac{n}{2}\left((x+y)\left(\frac{1}{2}+\eta\right)-y\right)^{2}\!+\frac{n}{2}\left((x+y)\left(\frac{1}{2}-\eta\right)-x\right)^{2}$
@@
Let $\epsilon=\|\Pi P_{\theta}+P_{\theta}\Pi-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}/(n\|\pi\|_{\infty})$ to ensure that the $P_{\theta}$’s satisfy the condition of the alternative hypothesis in (17). Substituting the values of $\|\pi\|_{\infty}=y$ and $\|\Pi P_{\theta}+P_{\theta}\Pi-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\mathrm{F}}$ implies that
$\epsilon^{2}\leq\frac{4\eta^{2}}{n}.$ (88)
Now, the Ingster-Suslina method [IngsterSuslina2003] states that
$\mathcal{R}_{\mathsf{m}}[{\mathcal{G}},\epsilon]\geq 1-\sqrt{\frac{\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})}{2}},$ (89)
where $\chi^{2}(\cdot||\cdot)$ denotes the $\chi^{2}$-divergence. We compute $\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ by expressing it as an expectation with respect to two independent pairwise models corresponding to permutations $\theta$ and $\theta^{\prime}$ drawn independently and uniformly at random from $\Theta$ as
$1+\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ $={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\left[\int\frac{d{\mathbb{P}}_{\theta}d{\mathbb{P}}_{\theta^{\prime}}}{d{\mathbb{P}}_{0}}\right],$
where $d{\mathbb{P}}_{\theta}$ denotes the measure induced by the pairwise comparison model corresponding to permutation $\theta$. Let $p_{ij}$ and $p_{ij}^{\prime}$ be the pairwise probabilities corresponding to permutations $\theta$ and $\theta^{\prime}$. Now we will essentially utilize the tensorization property of $1+\chi^{2}(P||Q)$, i.e., we utilize the fact that for distributions $P_{1},Q_{1},\dots,P_{n},Q_{n}$, we have
@@
Therefore, $\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ can be simplified as
$1+\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\left[\prod_{i=1}^{n}\prod_{\begin{subarray}{c}j=1:\\ j\neq i\end{subarray}}^{n}\right$.
[…]
$={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\left[\prod_{i=1}^{n}\prod_{\begin{subarray}{c}j=1:\\ j\neq i\end{subarray}}^{n}\right$.
[…]
This calculation of simplifying the $\chi^{2}$-divergence follows an approach similar to that in [RastogiBalakrishnanShahAarti2022], but we include the details here for completeness. We will focus on the $(i,j)$th term of the product in Equation 90 for $i\neq j$ and denote it as $f(p_{ij},p^{\prime}_{ij})$:
@@
By our construction of pairwise comparison matrices both $p_{ij},p^{\prime}_{ij}$ take values in set $\{\frac{1}{2},\frac{1}{2}+\eta\}$ if $j\geq i$ (and $\{\frac{1}{2},\frac{1}{2}-\eta\}$ otherwise). Furthermore, whenever either $p_{ij}$ or $p_{ij}^{\prime}$ equals $\frac{1}{2}$, we have $f(p_{ij},p^{\prime}_{ij})=1$. Additionally, by Eq. 91, we have $f(\frac{1}{2}-\eta,\frac{1}{2}-\eta)=f(\frac{1}{2}+\eta,\frac{1}{2}+\eta)$. Let a random variable $B$ denote the number of occurrences where $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}+\eta$ in randomly drawn permutations $\theta$ and $\theta^{\prime}$ (or equivalently, $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}-\eta$). Consequently, we obtain
@@
This is because, by translated skew-symmetry, the number of occurrences of $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}+\eta$ is the same as the occurrences where $p_{ij}=p_{ij}^{\prime}=\frac{1}{2}-\eta$ and $f(p_{ij},p_{ij}^{\prime})$ is same in both the scenarios. Moreover, we can simplify $f(p_{ij},p_{ij}^{\prime})$ as
$f\left(\frac{1}{2}+\eta,\frac{1}{2}+\eta\right)$
$=2^{k}\sum_{m=0}^{k}{k\choose m}\left(\frac{1}{2}+\eta\right)^{2m}\left(\frac{1}{2}-\eta\right)^{2k-2m}$
@@
$=2^{k}\left(\frac{1}{2}+2\eta^{2}\right)^{k}\times$
$\ \ \ \ \ \ \ \left(\sum_{m=0}^{k}{k\choose m}\left(\frac{1}{2}+\frac{\eta}{\frac{1}{2}+2\eta^{2}}\right)^{m}\left(\frac{1}{2}-\frac{\eta}{\frac{1}{2}+2\eta^{2}}\right)^{k-m}\right)$
@@
$1+\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ $={\mathbb{E}}_{\theta,\theta^{\prime}\sim\text{Unif}(\Theta)}\bigg[(1+4\eta^{2})^{2kB}\bigg]$
$=\sum_{b=0}^{n/2}{\mathbb{P}}(B=b)(1+4\eta^{2})^{2kb}.$ (93)
To derive an upper bound on ${\mathbb{P}}(B=b)$, for a fixed permutation $\theta$, we aim to find another permutation $\theta^{\prime}$ with exactly $b$ aligned elements. These $b$ matches can be selected in ${n/2}\choose b$} ways. For the remaining n/2bn/2-b elements, we require derangements to avoid alignment with the perturbed elements in $\theta$. Therefore, the number of such permutations is $(\frac{n}{2}-b)!\sum_{i=0}^{n/2-b}\frac{(-1)^{i}}{i!}$. Clearly, this quantity is upper bounded by $\frac{1}{2}(\frac{n}{2}-b)!$, yielding an upper bound on ${\mathbb{P}}(B=b)$ as
${\mathbb{P}}(B=b)\leq\frac{\frac{1}{2}{{n/2}\choose b}(\frac{n}{2}-b)!}{(n/2)!}\leq\frac{1}{2(b!)}.$
@@
$\chi^{2}({\mathbb{P}}_{\Theta}||{\mathbb{P}}_{0})$ $\leq\sum_{b=0}^{n/2}\frac{1}{2(b!)}\bigg((1+4\eta^{2})^{2kb}-1\bigg)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ +\sum_{b=0}^{n/2}{\mathbb{P}}(B=b)-1$